OpenStudy (anonymous):

one more vector questions, please help |a|=2|b| (a+b) is vertical to (2a-5b) ask for the angle between a and b (a and b are not 0)

OpenStudy (anonymous):

@ZeHanz can you help again

OpenStudy (harsimran_hs4):

for angle between a and b you can use cosine formula (do you know it?) now to fill in the missing quantities in formula use (a+b) . (2a-5b) = 0 (because they are vertical/perpendicular) try it if you are struck we`ll analyze in detail

OpenStudy (harsimran_hs4):

*addition you don`t need cosine formula for this just start with the second hint

OpenStudy (anonymous):

ok i'll try, please give me a moment thank you very much

OpenStudy (anonymous):

oh ok i forgot what way did i use before but i'll think it over with your hint again

OpenStudy (harsimran_hs4):

use what ? well just go step by step with second hint and you`ll be trough in minutes

OpenStudy (anonymous):

ok, i know the formula of cos and i also know they are vertical so they dot and get 0 but what can i do with 2a^2=5b^2 sorry i don't know where I am stuck

OpenStudy (harsimran_hs4):

ok lets try step by step check this out (a+b) . (2a-5b) = 0 2a^2 - 5a.b - 2a.b - 5b^2 = 7a.b 3b^2 = 7ab cos0 ( here a and b on right side is magnitude of vector) clear till this step?

OpenStudy (harsimran_hs4):

*\[\cos \theta\] instead of cos 0

OpenStudy (anonymous):

sorry, why 2a^2-3ab-5b^2=7ab

OpenStudy (anonymous):

oh never mind, thank you very much, i know what had happened.

OpenStudy (anonymous):

i got it

OpenStudy (harsimran_hs4):

good !! (y)

OpenStudy (harsimran_hs4):

altough 2nd part 2a^2 - 5a.b - 2a.b - 5b^2 = 7a.b should have been 2a^2 - 5a.b - 2a.b - 5b^2 = 0